SUB-DOPPLER SLIT JET DISCHARGE SPECTROSCOPY OF JET COOLED POLYACETYLENES: THE ANTI-SYMMETRIC CH STRETCH MODE OF TRIACETYLENE

\underline{CHIH-HSUAN CHANG}, MELANIE A. ROBERTS, and DAVID J. NESBITT, JILA, National Institute of Standards and Technology University of Colorado, and Department of Chemistry and Biochemistry University of Colorado at Boulder, Colorado 80309.

Growth of polyacetylenic molecules in acetylene flames is thought to play a central role in combustion chemistry and formation of soot, as well as the chemistry of gas clouds in the interstellar medium. In this talk, we present results from first sub-Doppler, high resolution infrared spectroscopic studies on triacetylene.^{a}^{,}^{b} In particular, we explore the fundamental anti-symmetric CH stretching mode (v_{5}) of jet-cooled triacetylene in a pulsed slit discharge, where the in − situ synthesis arises from a discharge of trace(0.1-1%) acetylene/rare gas mixtures followed by rapid CCH + HCCH chemistry in the supersonic expansion environment. The band origin of this mode is determined to be 3329.0544(2) cm^{−1}. At high resolution, a series of avoided energy level crossings arising from rotational perturbations are observed and ascribed to perpendicular Coriolis mixing with a near degenerate manifold of Π vibrational symmetry. The energy level patterns are successfully analyzed to reveal spectroscopic constants and Coriolis coupling matrix elements for the perturbing manifold. In addition, a weak Π-Π hot band progression due to thermal population in the slit jet is observed and assigned.

aD. McNaughton and D. N. Bruget, J. Mol. Spectrosc.150, 620 (1991)

bK. Matsumura, K. Kawaguchi, D. McNaughton, and D. N. Bruget, J. Mol. Spectrosc.158, 489 (1993)